1. In each of the following piece-wise functions: (i) Sketch the graph of the given function, (ii) Express f(t) in terms of t Help, please.
Answer #1

- >ه Note : (1) a {ucty = (II) a {uct-01 (III) IF G (3) = L { get ak get-es) hen G() olt-e). é esoct<2 (a) f(t) = -(1-2) e t.7, 2 (i) then the groaph of fot) is given by24 0 5 10 15 -2Νσω + e-6-2) &{{e (t-2 _ Just-235 (W) f(t) = 1. 1. Zuct ult-0) - uct-2) ult-2) = u(t) + [4-(4-2) 1] uct-2) <{f(t)} =a1 <{ucts(b) t oct<2 f(t) = 2 9 7-t 2 <tL5 54 tl t77 0 (i) Then the is groach of the function f ct5. 0 5 10 15 - 5Νσω f(t) = t[ult-o) - ult-2)] + 2[uct-2) - u(t-1)] + (7-D) [ ult-5) - ult -7)] + 0. uct-7) tult) & (2-t) ult-2) + (5-t) ult-5のんけ 11 -00人イ竹 ちるんで+0 ねんで する -22 e e55 る 75 Q 2 32 g2 -25 e -5 75 e te